Suppose that you open your own business and eam an accounting proft of $40.00 pr year When you started your business you left a job that paid you a salary annually. profit is $25,000 Also, suppose that you invested $70,000 of your own funds to start up your business. If the normal rate of return on capital is 5 percent, your econam OA.-$11,500 OB. $11,500. OC.-$55,000. OD. $15,000.
Answer #1

Answer

The correct answer is (B) $11500

Economic Profit = Accounting Profit - Implicit Cost.

Implicit Cost can be the income you sacrificed because of starting a new Business, Interest or Return You foregone because of investing in the business.

Here, income you sacrificed because of starting a new Business = $25000 and

Interest or Return You foregone because of investing in the business = 5% of 70,000 = $3500

Hence Implicit Cost = 25000 + 3500 = $28500

It is also given that Accounting Profit = $40000

Hence,

Economic Profit = 40000 - 28500 = $11500

Hence, the correct answer is (B) $11500